In an investigation into a patient’s red corpuscle count, the number of such corpuscles
appearing in each of 400 cells of a haemocytometer was counted. The results were as
follows:

No. of red blood corpuscles 0 1 2 3 4 5 6 7 8
No. of cells 40 66 93 94 62 25 14 5 1

It is thought that a Poisson distribution with mean .. provides an appropriate model for this
situation.
(a) Estimate mu, the Poisson parameter.
(b) Test the fit of the Poisson model.
Answer #1

(a) Estimate of \mu

\bar{X}=\sum f.x / \sum f

= 2.585

(b)

The null hypothesis

H0: The data fits the Poisson model.

Test statistic

(Oi - Ei)2

where Oi : observed frequency

Ei : expected frequency

x Oi Ei (Oi-Ei)^2/Ei
0 40 30.15843 3.2115882
1 66 77.95955 1.83467907
2 93 100.7627 0.598035923
3 94 86.82387 0.593118212
4 62 56.10993 0.618303425
5 25 29.00883 0.553994587
6 14 12.49797 0.180516362
7 5 4.615322 0.032062071
8 1 1.491326 0.161870242
399.4279 7.784168093

Therefore,

\chi ^{2}=7.784

Note :Expected frequencies are calculated as , Ei = 400* P(X=x) = 400*\frac{e^{-\mu }\mu^{x}}{x!}

df = 8

P value = 0.4548

Since P value > 0.05

, result is not significant

Fail to reject H0.

There is not enough evidence to conclude that the data do not fit Poisson model.